OpenStudy (anonymous):

SIMPLIFY! :) y/3 + y/7 + y/5

OpenStudy (anonymous):

Create a common denominator, in this case it can be 105. \[\frac{ 35y + 15y + 21y }{ 105 }= \frac{ 71 y}{ 105 }\]

OpenStudy (anonymous):

Thanks!! !:)